# Stochastic Cahn–Hilliard Equations from One-Dimensional Ising–Kac–Kawasaki Dynamics

### 31 Mar 2026 v1

@@ 3 Expansion for the discrete drift
$|\mathcal{J}_{\gamma}^{\mathrm{rem}}(x)|\lesssim\frac{\varepsilon_{\gamma}^{4}\delta_{\gamma}}{\alpha_{\gamma}}\,\bigl|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(x)\bigr|^{2}.$
In particular, (14) makes transparent why the nonlinear drift closure is governed by the second-order Boltzmann–Gibbs replacement for $d_{x}^{2}$ and by the mesoscopic replacements needed to identify $(m^{\ell})^{2}$ with $\delta_{\gamma}^{2}X_{\gamma}^{2}$: after inserting the BG closure into $\mathcal{J}_{\gamma}^{\mathrm{nl}}$ and taking the outer discrete divergence, one obtains the cubic conservative term $\partial_{x}^{2}(X^{3})$ in the scaling limit.
For each $\phi\in C^{\infty}(\mathbb{T})$, we introduce the discrete Kac-smoothed test function
@@
Then the drift term can be separated as $\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}=\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)+\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi)+\mathcal{D}_{\gamma}^{\mathrm{rem}}(\phi),$, where
$\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi):=\varepsilon_{\gamma}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}\phi(x)\mathcal{J}_{\gamma}^{\mathrm{lin}}(t,x)=-\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}d_{x}\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}x)+o_{\gamma}(1),$
$\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi):=\varepsilon_{\gamma}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}\phi(x)\mathcal{J}_{\gamma}^{\mathrm{nl}}(t,x)=-\frac{\beta}{4}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}d_{x}^{2}\,\nabla_{\varepsilon_{\gamma}}X_{\gamma}(\varepsilon_{\gamma}x)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}x)+o_{\gamma}(1),$
$\mathcal{D}_{\gamma}^{\mathrm{rem}}(\phi):=\varepsilon_{\gamma}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}\phi(x)\mathcal{J}_{\gamma}^{\mathrm{rem}}(\varepsilon_{\gamma}x)=o_{\gamma}(1).$
@@
$\int_{0}^{t}\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)(\sigma\big(\frac{s}{\alpha_{\gamma}}\big))\,ds=-A^{\prime}_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}\,ds-\nu_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}\,ds+o_{\gamma}(1)$
@@
$\nu_{\gamma}:=\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\mathfrak{m}_{2,\gamma}}{2}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\gamma^{2}}\to\frac{\lambda}{4}\mathfrak{m}_{2}=\nu,\quad\text{and}\quad A^{\prime}_{\gamma}:=\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\to\frac{1}{2}\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}:=A^{\prime}.$
@@
Recall that the linear part is defined as
$\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi):=-\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}d_{x}\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(x)+o_{\gamma}(1)$ (16)
Note that $d_{x}=\sigma_{x}-\sigma_{x+1}$ is the conservative gradient of the spin field, and $X_{\gamma}$ is defined by the coarse-grained and rescale of the spin field. Using the discrete integration by parts, the linear part $\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)$ can be rewritten as
@@
$\int_{0}^{t}\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)(\sigma\big(\frac{t}{\alpha_{\gamma}}\big))\,ds=-A^{\prime}_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}\,ds-\nu_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}\,ds+o_{\gamma}(1)$
@@
By Theorem 3, the rescaled field $X_{\gamma}$ satisfies the following discrete SPDE on lattice $\Lambda_{N}$:
@@
In order to obtain the uniform moment bound and the uniform drift bound needed in the tightness argument, it is more convenient to work with the Da Prato–Debussche type decomposition of the rescaled field rather than to apply Itô’s formula directly to the full process $X_{\gamma}$.
By using the discrete semigroup $S_{\gamma}:=e^{-\nu_{\gamma}t\Delta^{2}_{\varepsilon_{\gamma}}}$, we define the discrete stochastic convolution $Z_{\gamma}:=\int^{t}_{0}S_{\gamma}(t-s)M_{\gamma}(t)$, so that
@@
$\partial_{t}Y_{\gamma}=-\nu_{\gamma}\Delta_{\varepsilon_{\gamma}}^{2}Y_{\gamma}-A^{\prime}_{\gamma}Y_{\gamma}+\mathcal{N}_{\gamma}(X_{\gamma})+\mathcal{R}_{\gamma}(X_{\gamma}),\qquad Y_{\gamma}(0)=X_{\gamma}(0).$ (22)
@@ 4 The Second-order Boltzmann–Gibbs Principle
Throughout this section, all limits are taken along a sequence $\gamma\downarrow 0,N=N_{\gamma}\to\infty.$ We denote by $\mu_{N,\gamma,\beta}$ the corresponding reversible Gibbs measure of the one-dimensional Ising–Kac–Kawasaki dynamics $\sigma_{t}$.
We assume that the initial law is $f_{N,0}\mu_{N,\gamma,\beta}$, where
$H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})\leq CN$
for some constant $C>0$, and that the initial empirical magnetization field converges to a macroscopic profile $m_{0}$.
Fix $r\geq 1$, and let $\Psi:\{-1,1\}^{\Lambda_{r}}\to\mathbb{R}$ be a bounded local observable. For $\ell>r$, define the centered block
@@
and the corresponding block average
@@
With $L\gg\ell$, we also define the mesoscopic block average in two block estimates
@@
For each bond $z=(z,z+1)\in\Lambda_{N}$, define $A_{z}(\sigma):=\sigma(z+1)-\sigma(z),$ and let
$D_{z}^{N,\gamma}(f):=\frac{1}{4}\int c_{z}(\sigma)\,\bigl(f(\sigma^{z,z+1})-f(\sigma)\bigr)^{2}\,\mu_{N,\gamma,\beta}(d\sigma)$
be the local Dirichlet form associated with the Kawasaki exchange on the bond $(z,z+1)$.
In this section, we also assume that following assumption hold.
Assumption 4 (Block scales).
There exist mesoscopic block sizes $\ell=\ell_{\gamma},L:=L_{\gamma}\in\mathbb{N}$ such that
$\ell=\ell(\gamma)\to\infty,\quad L=L(\gamma)\to\infty,\quad\frac{\ell}{L}\to 0,\quad\frac{L}{N}\to 0,\quad\gamma\ell\to 0,\quad\gamma^{2}L^{3}\to 0.$
Whenever the weighted block-to-Kac replacement is used, we further assume
$\sup_{\gamma\downarrow 0}\frac{\varepsilon_{\gamma}^{2}\gamma^{1/2}}{\alpha_{\gamma}\delta_{\gamma}}<\infty.$

### 05 Apr 2026 v2

@@ 3 Expansion for the discrete drift
$\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}k)=h_{\gamma}\!\left(\sigma^{N}\left(\frac{t}{\alpha_{\gamma}}\right);k+1\right)-h_{\gamma}\!\left(\sigma^{N}\left(\frac{t}{\alpha_{\gamma}}\right);k\right).$
Hence
$\bigl|\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}k)\bigr|$ $=\left|\sum_{j\in\Lambda_{N}}\bigl[\kappa_{\gamma}(k+1-j)-\kappa_{\gamma}(k-j)\bigr]\sigma_{j}\!\left(\frac{t}{\alpha_{\gamma}}\right)\right|$
$\leq\sum_{j\in\Lambda_{N}}\bigl|\kappa_{\gamma}(k+1-j)-\kappa_{\gamma}(k-j)\bigr|,$
since $|\sigma_{j}|=1$. By translation invariance of the torus, the right-hand side does not depend on $k$, so
$\bigl|\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}k)\bigr|\leq\sum_{m\in\mathbb{Z}}|\kappa_{\gamma}(m+1)-\kappa_{\gamma}(m)|.$
It therefore remains to estimate the discrete variation of $\kappa_{\gamma}$. Using the Kac form,
$|\kappa_{\gamma}(m+1)-\kappa_{\gamma}(m)|\leq\,\gamma^{2}\int_{0}^{1}|\mathfrak{K}^{\prime}(\gamma m+\theta\gamma)|\,d\theta.$
Summing over $m\in\mathbb{Z}$, and $\mathfrak{K}\in W^{1,1}(\mathbb{R})$, we obtain
$\sum_{m\in\mathbb{Z}}|\kappa_{\gamma}(m+1)-\kappa_{\gamma}(m)|\leq C\gamma\|\mathfrak{K}\|_{W^{1,1}(\mathbb{R})}.$
Therefore
$\sup_{0\leq t\leq T}\sup_{x\in\Lambda_{\varepsilon_{\gamma}}}\bigl|\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,x)\bigr|\leq C\gamma,$
which proves the claim. ∎
Now we show the remainder term satisfies uniform estimate $|\nabla_{\varepsilon_{\gamma}}\mathcal{J}_{\gamma}^{\mathrm{rem}}|=o_{\gamma}(1)$.
Theorem 3.
Assume that $\frac{\gamma^{4}}{\alpha_{\gamma}\delta_{\gamma}}\rightarrow 0$ as $\gamma\downarrow 0$. Then for every $T>0$,
$\sup_{0\leq t\leq T}\sup_{x\in\Lambda_{\varepsilon_{\gamma}}}\bigl|\nabla_{\varepsilon_{\gamma}}\mathcal{J}_{\gamma}^{\mathrm{rem}}\bigr|\rightarrow 0\quad\text{as}\quad\gamma\downarrow 0.$
Proof.
Writing $x=\varepsilon_{\gamma}k$, we get
$\nabla_{\varepsilon_{\gamma}}\mathcal{J}_{\gamma}^{\mathrm{rem}}(t,x)=-\frac{1}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\bigl[\kappa_{\gamma}(k+1-i)-\kappa_{\gamma}(k-i)\bigr]j_{i}^{\mathrm{rem}}(t),$
Since $F(r)=\frac{1}{1+e^{r}}$ is smooth and satisfies $F^{\prime\prime}(z)\to 0,$ as $z\to 0$, the mean value theorem yields
$|F^{\prime\prime}(a)|\leq\|F^{(3)}\|_{L^{\infty}(\mathbb{R})}\,|a|\qquad\text{for all }a\in\mathbb{R}.$
Hence $|F^{\prime\prime}(a_{i})|\leq C\,\kappa_{\gamma}(1).$ Moreover, Lemma 10 yields that
$|b_{i}|=\bigl|\beta d_{i}\,\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,x_{i})\bigr|\leq C\gamma\beta,$
uniformly for $0\leq t\leq T$. Therefore $|j_{i}^{(2),\gamma}(t)|\leq C\,\kappa_{\gamma}(1)\gamma^{2}$ uniformly in $i,t,\gamma$.
By Taylor’s theorem with Lagrange remainder, there exists $\theta_{i}^{\gamma}\in(0,1)$ such that
$\widetilde{R}_{i}^{\gamma}=\frac{1}{6}\,d_{i}\,F^{(3)}(a_{i}+\theta_{i}^{\gamma}b_{i})\,(b_{i})^{3}.$
Since $F^{(3)}$ is bounded and $|d_{i}|\leq 2$, Lemma 10 yields $|\widetilde{R}_{i}^{\gamma}|\leq C\,|b_{i}|^{3}\leq C\gamma^{3}.$ Since $\mathfrak{K}\in W^{1,1}(\mathbb{R})$ and $\kappa_{\gamma}(z)=\gamma\mathfrak{K}(\gamma z)$, a discrete mean-value estimate yields
$\sum_{j\in\mathbb{Z}}|\kappa_{\gamma}(j+1)-\kappa_{\gamma}(j)|\leq C\gamma.$
Thus
$\sup_{0\leq t\leq T}\sup_{x\in\Lambda_{\varepsilon_{\gamma}}}\bigl|\nabla_{\varepsilon_{\gamma}}\mathcal{J}_{\gamma}^{\mathrm{rem}}(t,x)\bigr|$
$\leq$ $\frac{1}{\alpha_{\gamma}\delta_{\gamma}}\sup_{i}|j_{i}^{(2),\gamma}(t)|\sum_{j\in\mathbb{Z}}|\kappa_{\gamma}(j+1)-\kappa_{\gamma}(j)|+\frac{1}{\alpha_{\gamma}\delta_{\gamma}}\sup_{i}|\widetilde{R}_{i}^{\gamma}(t)|\sum_{j\in\mathbb{Z}}|\kappa_{\gamma}(j+1)-\kappa_{\gamma}(j)|$
$\leq$ $C\,\frac{\gamma^{3}\kappa_{\gamma}(1)}{\alpha_{\gamma}\delta_{\gamma}}+C\,\frac{\gamma^{4}}{\alpha_{\gamma}\delta_{\gamma}}\to 0$
This proves the claim. ∎
For every test function $\phi\in C^{\infty}(\mathbb{T})$, we define the weak drift components by
$\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi;\sigma)$ $:=\bigl\langle\mathcal{J}_{\gamma}^{\mathrm{lin}}(\sigma),\nabla_{\varepsilon_{\gamma}}\phi\bigr\rangle_{\gamma}=\varepsilon_{\gamma}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}\mathcal{J}_{\gamma}^{\mathrm{lin}}(\sigma;x)\,\nabla_{\varepsilon_{\gamma}}\phi(x),$
$\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;\sigma)$ $:=\bigl\langle\mathcal{J}_{\gamma}^{\mathrm{nl}}(\sigma),\nabla_{\varepsilon_{\gamma}}\phi\bigr\rangle_{\gamma}=\varepsilon_{\gamma}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}\mathcal{J}_{\gamma}^{\mathrm{nl}}(\sigma;x)\,\nabla_{\varepsilon_{\gamma}}\phi(x),$
$\mathcal{D}_{\gamma}^{\mathrm{rem}}(\phi;\sigma)$ $:=\bigl\langle\mathcal{J}_{\gamma}^{\mathrm{rem}}(\sigma),\nabla_{\varepsilon_{\gamma}}\phi\bigr\rangle_{\gamma}=\varepsilon_{\gamma}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}\mathcal{J}_{\gamma}^{\mathrm{rem}}(\sigma;x)\,\nabla_{\varepsilon_{\gamma}}\phi(x).$
Therefore,
$\Bigl(\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}\Bigr)(\sigma)=\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi;\sigma)+\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;\sigma)+\mathcal{D}_{\gamma}^{\mathrm{rem}}(\phi;\sigma).$
Now we introduce the discrete Kac-smoothed test function
@@
Then the linear term
$\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi;\sigma)$ $=\varepsilon_{\gamma}\sum_{x}\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)d_{i}(\sigma)\nabla_{\varepsilon_{\gamma}}\phi(x)$
$=\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}d_{i}(\sigma)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)$
$=-\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}\sigma_{i}\,\Delta_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i).$ (19)
Similarly, the nonlinear term
$\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;\sigma)=\frac{\beta\varepsilon_{\gamma}^{3}}{4\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}d_{i}(\sigma)^{2}\,\nabla_{\varepsilon_{\gamma}}X_{\gamma}(\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i).$
Moreover, the remainder term $\mathcal{D}_{\gamma}^{\mathrm{rem}}(\phi)=o_{\gamma}(1).$
Now we consider the expansion for the linear part $\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)$.
@@
$\int_{0}^{t}\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)\Big(\sigma^{N}\big(\frac{s}{\alpha_{\gamma}}\big)\Big)\,ds=-A^{\prime}_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}\,ds-\nu_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}\,ds+o_{\gamma}(1)$
@@
$\nu_{\gamma}:=\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\mathfrak{m}_{2,\gamma}}{2}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\gamma^{2}}\to\frac{\lambda}{4}\mathfrak{m}_{2}=\nu,\quad\text{and}\quad A^{\prime}_{\gamma}:=\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\to\frac{1}{2}\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}:=\frac{\lambda}{2}.$
@@
Using the discrete integration by parts, the linear part $\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)$ defined in (3) can be rewritten as
@@
$\int_{0}^{t}\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)\Big(\sigma^{N}\big(\frac{t}{\alpha_{\gamma}}\big)\Big)ds=-A^{\prime}_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}\,ds-\nu_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}\,ds+o_{\gamma}(1)$
@@
Combining with Theorem 3 and Theorem 4, the rescaled field $X_{\gamma}$ satisfies the following discrete SPDE on lattice $\Lambda_{\varepsilon_{\gamma}}$:
@@
Moreover, the fast process $d_{i}^{2}(t/\alpha_{\gamma})=(\sigma_{i+1}(t/\alpha_{\gamma})-\sigma_{i}(t/\alpha_{\gamma}))^{2}$ in (16) implies that the nonlinear drift $\mathcal{N}(X_{\gamma})$ is governed by the second-order Boltzmann–Gibbs replacement $\delta_{\gamma}^{2}X_{\gamma}^{2}$ for $d_{x}^{2}$.
In order to obtain the uniform moment bound and the uniform drift bound needed in the tightness argument, it is more convenient to work with the Da Prato–Debussche type decomposition of the rescaled field rather than to apply Itô’s formula directly to the full process $X_{\gamma}$. By using the discrete semigroup $S_{\gamma}:=e^{-\nu_{\gamma}t\Delta^{2}_{\varepsilon_{\gamma}}}$, we define the discrete stochastic convolution $Z_{\gamma}:=\int^{t}_{0}S_{\gamma}(t-s)M_{\gamma}(t)$, so that
@@
$\partial_{t}Y_{\gamma}=-\nu_{\gamma}\Delta_{\varepsilon_{\gamma}}^{2}Y_{\gamma}-A^{\prime}_{\gamma}\Delta_{\varepsilon_{\gamma}}Y_{\gamma}+\mathcal{N}_{\gamma}(X_{\gamma})+\mathcal{R}_{\gamma}(X_{\gamma}),\qquad Y_{\gamma}(0)=X_{\gamma}(0).$ (25)
@@ 4 The Second-order Boltzmann–Gibbs Principle
In this section, we establish the second-order Boltzmann–Gibbs principle for the fast Kawasaki dynamics and use it to close the nonlinear conservative drift identified in Section 3. More precisely, recall that the nonlinear part of the weak drift is written in current form as $\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;t)=\bigl\langle\mathcal{J}_{\gamma}^{\mathrm{nl}}(t),\nabla_{\varepsilon_{\gamma}}\phi\bigr\rangle_{\gamma},$ so that the key microscopic quantity to be replaced is the local bond observable $d_{i}^{2}(t):=\bigl(\sigma_{i}(t/\alpha_{\gamma})-\sigma_{i+1}(t/\alpha_{\gamma})\bigr)^{2}.$ The purpose of this section is to show that, after a conservative multiscale replacement, $d_{i}^{2}$ can be approximated by an explicit local function of the coarse field. Combined with the current structure from Section 3, this yields the cubic conservative term $\Delta(X^{3})$ in the limiting stochastic Cahn–Hilliard equation.
The argument proceeds in three steps. First, a one-block estimate replaces a local observable by its conditional expectation under a local canonical Gibbs state with the same block magnetization. Second, a two-block estimate allows one to compare neighboring block averages and to pass from microscopic to mesoscopic block scales. Third, a Kac-scale matching argument identifies the resulting mesoscopic quantity with a function of the Kac coarse-grained field $X_{\gamma}$. This produces the required closure of the nonlinear weak drift.
Throughout this section, all limits are taken along a sequence $\gamma\downarrow 0,N=N_{\gamma}\to\infty.$ We work on the canonical sector associated with the fixed total magnetization, and denote by $\mu_{N,\gamma,\beta}$ the corresponding reversible Gibbs measure for the one-dimensional Ising–Kac–Kawasaki dynamics.
Fix $r\geq 1$, and let $\Psi:\{-1,1\}^{\Lambda_{r}}\to\mathbb{R}$ be a bounded local observable. For $\ell>r$, define
@@
and the associated block average
@@
For a larger block scale $L\gg\ell$, we also define
@@
Since the nonlinear current is built from bond variables, it is convenient to introduce
$d_{z}(\sigma):=\sigma_{z}-\sigma_{z+1},\qquad z\in\Lambda_{N},$
and, for every density $f$ on the canonical sector, the local Dirichlet form
$D_{z}^{N,\gamma}(f):=\frac{1}{4}\int c_{z}(\sigma)\,\bigl(f(\sigma^{z,z+1})-f(\sigma)\bigr)^{2}\,\mu_{N,\gamma,\beta}(d\sigma),$
associated with the Kawasaki exchange across the bond $(z,z+1)$. These local Dirichlet forms will be the basic input in the entropy–Kipnis–Varadhan bounds used below.
We now impose the block-scale assumptions under which the one-block and two-block replacement estimates are carried out.
Assumptions for block scales.
Throughout this section, we fix two mesoscopic block scales
$\ell_{\gamma},\;L_{\gamma}\in\mathbb{N},\qquad\ell_{\gamma}<L_{\gamma},$
such that, as $\gamma\downarrow 0$,
$\ell_{\gamma}\to\infty,\qquad L_{\gamma}\to\infty,\qquad\frac{\ell_{\gamma}}{L_{\gamma}}\to 0,\qquad\frac{L_{\gamma}}{N}\to 0,\qquad\frac{\gamma^{2}L_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}\to 0\qquad\gamma^{2}L_{\gamma}^{3}\to 0.$
These conditions are the basic scale assumptions for the one-block and two-block replacements.
Under these scale assumptions, we first replace a local observable by its canonical Gibbs conditional expectation on a block of size $\ell_{\gamma}$, and then compare the resulting block average with a larger mesoscopic block average of size $L_{\gamma}$.